# Initialization using Update Approximation is a Silver Bullet for Extremely Efficient Low-Rank Fine-Tuning

### v1

@@ Appendix B Inference Costs
RoBERTa-large LoRA 88 $77.86$ G $155.79$ G
@@ Appendix C Experiment Details
Table 8: hyperparameter settings for RoBERTa-large on GLUE.
@@ Appendix D Dataset Details
The MetaMathQA dataset (Yu et al., 2024) creates mathematical questions by rephrasing existing ones from different viewpoints, without adding new information. We assess this dataset using two benchmarks: GSM8K (Cobbe et al., 2021), which consists of grade-school math problems requiring multi-step reasoning, and MATH (Hendrycks et al., 2021), which presents difficult, competition-level math problems.
CommonSense170K is a comprehensive dataset that consolidates eight commonsense reasoning datasets (Hu et al., 2023), as described below:

### v3

@@ Appendix B Inference Overhead vs LoRA
RoBERTa-large LoRA $8$ $77.86$ G $155.79$ G
@@ Appendix C Experiment Details
For all tasks, we followed the baseline configurations provided in the PiSSA (Meng et al., 2024) and rsLoRA (Kalajdzievski, 2023) papers for our comparisons.
@@
Table 8: Hyperparameter settings for RoBERTa-large on GLUE.
@@ Appendix D Dataset Details
The MetaMathQA dataset (Yu et al., 2024) creates mathematical questions by rephrasing existing ones from different viewpoints, without adding new information. We assess this dataset using two benchmarks: GSM8K (Cobbe et al., 2021), which consists of grade-school math problems requiring multi-step reasoning, and MATH (Hendrycks et al., 2021), which presents difficult, competition-level math problems. Evaluation focuses solely on the final numeric answer.
CommonSense170K is a comprehensive dataset that consolidates eight commonsense reasoning datasets (Hu et al., 2023). Each example is framed as a multiple-choice question where the model generates the correct answer without explanations. We use the prompt template from (Hu et al., 2023). The individual datasets used are described below: